# GlaGAN: A Generative Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

### v4

@@ 5.7 Analysis
Optimal selection of $R_{min}$ values allows GlaGAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1, Table 2, Table 4). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [29, 30, 31, 33, 32]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted GlaGAN to achieve amazing segmentation results. Unlike GAN+SLSMs or SLSMs, GlaGAN does not require additional data to avoid fitting issues. Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation.
To the best of our knowledge, GlaGAN is the best model currently, tailored to segment OCTA retinal main vessel.
@@ 6 Conclusion
In this paper, we propose GlaGAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation. Extensive experiments demonstrate that GlaGAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
Acknowledgment
This research is partially supported by NASA under grant 80NSSC22K0144, UTSW GMO No. 241213, and NIH grant 1R01AG083179-01.

### v5

@@ 5.5 Analysis
Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1 and Table 3). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [34, 35, 36, 37]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results. Unlike GAN+SLSMs or SLSMs, X-GAN does not require additional data to avoid fitting issues. Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation. Studies [32, 33] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models.
@@ 6 Conclusion
In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
7 Future Work
With access to state-of-the-art glaucoma detection equipment, we have curated a high-quality image dataset featuring precisely annotated retinal vascular images. This dataset is well-suited for glaucoma pathology research and computer vision.